OpenStudy (anonymous):

tyrone saved $24 of the $50 he earned mowing lawn. What percent of his earning did Tyrone Save?

OpenStudy (anonymous):

divide 24 by 50

OpenStudy (anonymous):

you get 24/50 = 12/25 = 6/12.5 = 3 / 6.25 = 1.5 / 3.125

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

= 0.48

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sorry now in percent form multiply by 100 so you get 48%

OpenStudy (anonymous):

no problem!